Consider the given right triangle with legs of length x cm. and y cm. and angle θ radians. If x is decreasing at the rate of 3 cm/min. and y is increasing at the rate of 4 cm/min., at what rate is angle θ changing when x=5 cm. and y=2 cm. ?


1
Expert's answer
2021-10-12T02:05:29-0400

Solution:

"dy/dt=4\ cm/min
\\ dx/dt=-3\ cm/min
\\x=5\ cm
\\y=2\ cm"

"\sec\theta=\dfrac {\sqrt{x^2+y^2}}{x}
\\\Rightarrow \sec^2\theta=\dfrac {{x^2+y^2}}{{x^2}} \ \ ...(i)"

Also,

"\tan \theta=\dfrac{AC}{AB}
\\\Rightarrow\tan \theta=\dfrac yx"

On differentiating both sides w.r.t "t",

"\sec^2\theta \dfrac{d\theta}{dt}=\dfrac{x(dy/dt)-y(dx/dt)}{x^2}
\\\Rightarrow\dfrac {x^2+y^2}{{x^2}} \dfrac{d\theta}{dt}=\dfrac{x(dy/dt)-y(dx/dt)}{x^2}
\\ \Rightarrow (\dfrac{5^2+2^2}{5^2}) \dfrac{d\theta}{dt}=\dfrac{5(4)-2(-3)}{5^2}"

"\\ \Rightarrow (\dfrac{29}{25}) \dfrac{d\theta}{dt}=\dfrac{26}{25}
\\ \Rightarrow \dfrac{d\theta}{dt}=\dfrac{26}{29} \ radians/min"